# The Power of Graph Sparsification in the Continual Release Model

### v1

@@ 8.1 Inner Product Queries
Theorem 8.1 (Theorem 4.5 in [jain2023counting]).
@@
8.2 Maximum Cardinality Matching
@@ 20       $r^{(0)},r^{(1)},\dots,r^{(k)}\leftarrow\mathcal{A}(S)$ answers to queries
Algorithm 8.1 Reduction from Inner Product Queries to Maximum Matching
Lemma 8.2.
Given an $\varepsilon$-DP mechanism for maximum matching in the continual release setting that outputs estimates within additive error at most $\zeta$ with probability $0.99$ for a fully-dynamic stream of length $T\geq n+2nk$, Algorithm 8.1 is an $\varepsilon$-DP mechanism for answering $k$ inner product queries within additive error $2\zeta$ with probability $0.99$.
Proof (Lemma 8.2).
Let $\mathcal{M}$ be such a $\varepsilon$-DP mechanism and suppose we are given a private dataset $y\in\{0,1\}^{n}$ as well as public linear queries $q^{(1)},\dots,q^{(k)}$. Let $S^{(0)},S^{(1)},\dots,S^{(k)}$ be the partial update streams as in Algorithm 8.1 and
@@
Proof (Theorem 8.3).
Let $c_{1},c_{2}$ be the constants from Theorem 8.1. Fix $n\in\mathbb{N}$ and suppose $T\geq n+2c_{1}n^{2}$. By Lemma 8.2, $\mathcal{A}$ implies a $1$-DP mechanism for answering $k=c_{1}n$ inner product queries within additive error $2\zeta$ with probability $0.99$. But then by Theorem 8.1, we must have $\zeta\geq c_{2}\sqrt{n}/2$.
@@
In the regime $T\geq\nicefrac{{1}}{{\varepsilon}}$, combining Lemma 8.4 and Theorem 8.3 implies a lower bound of
@@
8.3 Connected Components
@@ 24             Output $\lVert q^{(j)}\rVert_{0}+r^{{(j)}}_{n}-r^{(0)}_{3n}$
Algorithm 8.2 Reduction from Inner Product Queries to Connected Components
Lemma 8.6.
Given an $\varepsilon$-DP mechanism for connected components in the continual release setting that outputs estimates within additive error at most $\zeta$ with probability $0.99$ for a fully-dynamic stream of length $T\geq 3n+2nk$, Algorithm 8.2 is an $\varepsilon$-DP mechanism for answering $k$ inner product queries within additive error $2\zeta$ with probability $0.99$.
Proof (Lemma 8.6).
Let $\mathcal{M}$ be such a $\varepsilon$-DP mechanism and suppose we are given a private dataset $y\in\{0,1\}^{n}$ as well as public linear queries $q^{(1)},\dots,q^{(k)}$. Let $S^{(0)},S^{(1)},\dots,S^{(k)}$ be the partial update streams as in Algorithm 8.2 and
@@
Proof (Theorem 8.7).
Let $c_{1},c_{2}$ be the constants from Theorem 8.1. Fix $n\in\mathbb{N}$ and suppose $T\geq 3n+2c_{1}n^{2}$. By Lemma 8.6, $\mathcal{A}$ implies a $1$-DP mechanism for answering $k=c_{1}n$ inner product queries within additive error $2\zeta$ with probability $0.99$. But then by Theorem 8.1, we must have $\zeta\geq c_{2}\sqrt{n}/2$.
@@
By following an identical argument to Lemma 8.4, i.e. creating duplicate instances of the lower bound construction within a longer stream, we derive a similar lower bound to Theorem 8.5 that accounts for the privacy parameter $\varepsilon$.
Theorem 8.8.
8.4 Further Graph Statistics
We remark that the underlying idea for the basic maximum matching and connected components lower bounds (Theorem 8.3, Theorem 8.7) is that we can encode the bits of a secret database $y\in\{0,1\}^{n}$ within the structure of a sparse graph. Privately answering an inner product query on this database then reduces to answering a “bitwise OR” query by the inclusion-exclusion principle. The general bound which accounts for the privacy parameter $\varepsilon$ (Theorem 8.5, Theorem 8.8) is obtained through a folklore reduction to the basic lower bound.
It is not hard to see that this technique can extend to $k$-edge-connected component queries, $k$-vertex-connected component queries, and triangle counting queries for sparse graphs.
9 Conclusion & Future Work
In this paper, we initiated the study of low-space continual release algorithms for general graph problems. Using techniques from the non-private graph sparsification literature, we provided continual release algorithms for a variety of general graph problems, achieving, for the first time, nearly the same space and approximation guarantees of their non-private streaming counterparts. The improved space bounds are especially relevant for enabling computations on massive datasets, which are the core motivation of the field of online streaming algorithms. In addition to our space gains, we provide the first bounded arboricity sparsifiers for node-DP algorithms in the continual release model. We hope such sparsifiers have implications in the static DP model.
For our upper bounds, we mostly focused on the insertion-only setting of continual release. As the area of fully-dynamic algorithms in the continual release model is largely unexplored, we believe that an interesting future research direction is closing the gap in our theoretical understanding of this model. As observed in prior work, and hinted by our hardness results, the fully dynamic setting is significantly harder in continual release, with even basic graph problems requiring $\tilde{\Omega}(\operatorname*{poly}(n))$ additive error for dynamic streams while admitting $\tilde{O}(\operatorname*{poly}(\log(n))/\varepsilon)$ additive error in the insertion-only case. In this context, it would be especially interesting to deepen our understanding of the interplay between the dynamicity of the continual release setting (insertion-only vs fully-dynamic) and the space lower bounds (as opposed to error lower bounds) imposed by privacy. This is an area that has only recently received attention [dinur2023differential] and is an interesting future direction to explore.
[…]
Appendix A Related Works
Our work is related to multiple areas of algorithms including private graph algorithms, private streaming algorithms, and related non-private variants. We now cover the most relevant work in each area.
Streaming Algorithms
The streaming model of computation is a prominent model for large-scale data analysis that has been studied for multiple decades [morris1978counting]. In this model, one usually seeks space- and time-efficient algorithms that can process the stream on-the-fly without the need to store all data. A long line of work in this area includes classical results such as the Flajolet-Martin algorithm for counting distinct elements [flajolet1985probabilistic] and many others [flajolet2007hyperloglog, alon1996space]. In the context of streaming graph algorithms, ideally one would like to obtain space sublinear in the number of vertices [mcgregor2018simple, HuangP19]; but for many graph problems it is necessary to work in the semi-streaming model settling on space near-linear in the number of vertices [assadi2022semi, AG11, Esfandiari2018, feigenbaum2005graph, FMU22, mcgregor2015densest]. Online streaming algorithms release graph statistics after every update while maintaining the low space bounds [chen2023sublinear, halldorsson2016streaming, cormode2018approximating, GS24].  In the non-private streaming setting, there exists many works on graph algorithms including approximating the size of the maximum matching [AG11, ALT21, Assadi22, FMU22, mcgregor2018simple], vertex cover [assadi2019coresets], densest subgraph [EHW16, mcgregor2015densest], $k$-core decomposition [Esfandiari2018, KingTY23], number of connected components [BerenbrinkKM14], and others [assadi2022semi, AG11, feigenbaum2005graph, FMU22, HuangP19, henzinger1998computing, muthukrishnan2005data, mcgregor2014graph]
Continual Release Model
In the context of streaming computation, the DP model of reference is the continual release model [DworkNPR10, CSS11] where we require algorithms to abide by a strong privacy notion: an observer obtaining all future outputs of the algorithm must in essence learn almost nothing about the existence of any single input. Since its introduction, this research area has received vast attention outside of graphs, including many recent works (see e.g. [ChanLSX12, henzinger2024unifying, henzinger2023differentially, fichtenberger2023constant, jain2023counting, JRSS21]).
Among the insertion-only continual release work, prior research has tackled classical estimation problems [CSS11, henzinger2024unifying, henzinger2023differentially], as well as heavy hitters-related problems [ChanLSX12, epasto2023differentially]. More recent works tackle the fully-dynamic continual release setting [jain2023counting, dupre2023differential, FichtenbergerHO21]. Particularly relevant is [FichtenbergerHO21] that shows hardness results for graph estimation in fully-dynamic streams. Our work contributes new hardness results in this new emerging area as well.
Most relevant to our paper is the literature on continual release algorithms in graphs [upadhyay2021differentially, FichtenbergerHO21, Song18, jain2024time]. In this area, [Song18] studied graph statistics (degree distribution, subgraph counts, etc.) on bounded degree graphs. [FichtenbergerHO21] focused on estimating a number of graph statistics like maximum matching, triangle counting, and the density of the densest subgraph for both edge and node privacy; they provide approximation guarantees in terms of the maximum degree in the graph as well as lower bounds in insertion-only and fully dynamic streams. [jain2024time] focuses on counting problems in graphs (counting edges, triangles, stars, connected components), for node-privacy and where privacy must hold for arbitrary graphs (e.g., graphs with arbitrary degrees). They give time-aware projection algorithms that can transform any continual release algorithm that gives approximation guarantees for bounded degree graphs into a truly private algorithm on nearly bounded degree graphs. Like in [jain2024time] for our algorithms that assume a public bound, say on the stream arboricity, this bound affects only the approximation guarantees but not the privacy claims, hence our algorithms satisfy their notion of truly private [jain2024time].
Private Graph Algorithms
The private literature on graph algorithms includes a large body of work on static graph algorithms (see e.g. [AU19, blocki2022privately, dinitz2024tight, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, NRS07, RS16, raskhodnikova2016differentially, Upadhyay13] and references therein). Aside from the problems we study, work in this area includes results on preserving graph cuts [AU19], the stochastic block model recovery [chen2023private], graph clustering [bun2021differentially, imola2023differentially], and many other areas.
Our paper focuses on the pure, $\varepsilon$-DP setting. In this setting, there has been a variety of recent works that we study for static graphs. The densest subgraph (DSG) problem has been extensively studied in the non-private streaming context [BHNT15, EHW16, mcgregor2015densest]. Our work builds on the results of [EHW16, mcgregor2015densest] which show that edge sampling approximately preserves the density of the densest subgraph. In the context of privacy, beyond the already cited work of [FichtenbergerHO21] that focused on density estimation only, all other works are in the static DP or LEDP setting [NV21, AHS21, DLRSSY22, dinitz2024tight, DLL23]. In the $\varepsilon$-DP setting, the best known lower bound on the additive error of the densest subgraph problem is $\Omega(\sqrt{\varepsilon^{-1}\log n})$ [AHS21, NV21]. Comparatively, the best known upper bounds in the central $\varepsilon$-DP setting are the $(2,O(\varepsilon^{-1}\log n))$-approximation [DLL23], $(2+\eta,O(\varepsilon^{-1}\log^{2}n))$-approximate [dinitz2024tight], and the $(1+\eta,O(\varepsilon^{-1}\log^{4}n))$-approximation [DLRSSY22] upper bounds. In the $\varepsilon$-LEDP setting, the best known upper bounds are the $(2+\eta,O(\eta^{-1}\log^{2}n))$-approximate [dinitz2024tight] and $O(2,O(\varepsilon^{-1}\log n))$-approximate [DLL23] bounds. Our paper presents the first continual release algorithm for releasing an actual approximate densest subgraph (as opposed to estimating its density) in edge-private insertion-only streams.
To the best of our knowledge, the only differentially private algorithm that releases an implicit vertex cover is that of [gupta2010differentially] in the static $\varepsilon$-edge DP setting.
Appendix B The Size of a Maximum Matching
In this section, we recall the Micali-Vazirani algorithm [MV80] for finding a maximum matching in a graph. See also [vazirani2012simplification, vazirani1994theory] for a complete proof of correctness.
Theorem B.1 ([MV80]).
There is an algorithm that finds a maximum matching in a general graph $G=(V,E)$ in time $O(\lvert E\rvert\sqrt{\lvert V\rvert})$. Moreover, it uses $O(\lvert E\rvert)$ space.
The Micali-Vazirani algorithm repeatedly finds augmenting paths in each iteration. In each iteration, it finds a maximal set of disjoint minimum length augmenting paths with respect to the current matching and augments along all paths, both in $O(\lvert E\rvert)$ time. Then, one can argue that it suffices to terminate after $O(\sqrt{\lvert V\rvert})$ iterations [karzanov1973exact, HK71]. Each iteration can be implemented using $O(\lvert V\rvert+\lvert E\rvert)$ auxiliary space.
B.1 Coupled Node Sensitivity of the Greedy Matching Algorithm
Theorem B.2 (Section 1.4.2 in [blocki2022privately]).
The greedy matching algorithm $\textsc{GreedyMatching}_{\pi}$ that processes edges in the order specified by $\pi$ and outputs the size of a maximal matching has sensitivity 1 with respect to node neighboring graphs.
Appendix C Private Static Densest Subgraph
Theorem C.1 (Theorem 6.1 in [DLL23]).
@@ Appendix C Private Static Densest Subgraph
Theorem C.2 (Theorem 5.1 in [DLRSSY22]).
Fix $\eta\in(0,1]$. There is an $\varepsilon$-edge DP densest subgraph algorithm that runs in $O((m+n)\log^{3}n)$ time and $O(m+n)$ space and returns a subset $V^{\prime}\subseteq V$ of vertices that induces a $O\left(1+\eta,O(\varepsilon^{-1}\log^{4}n)\right)$-approximate densest subgraph with probability at least $1-O(1/\operatorname*{poly}(n))$ for any constant $\eta>0$.
Appendix D The Density of the Densest Subgraph
@@
Proposition D.2 ([chekuri2022densest]).
@@
Now we recall the classical result of [dinitz2006dinitz].
Theorem D.3 ([dinitz2006dinitz]).
@@
We are now ready to show Theorem D.1.
Proof (Theorem D.1).
@@
Construct the auxiliary flow network $D=(N,A)$ above for some guess $\lambda\geq\nicefrac{{\lvert E\rvert}}{{\lvert V\rvert}}$ of the value of the densest subgraph. There are $2+\lvert V\rvert+\lvert E\rvert=O(\lvert E\rvert)$ nodes and $\lvert E\rvert+2\lvert E\rvert+\lvert V\rvert=O(\lvert E\rvert)$ arcs. We can check if the densest subgraph has value at most $\lambda$ by computing the maximum flow in the $D$ using Dinic’s algorithm Theorem D.3. This is justified by Proposition D.2. Moreover, each flow computation also retrieves the vertices inducing a minimum $(s,t)$-cut by computing the vertices reachable from $s$ in the residual graph. The time and space complexity of computing the residual flow is dominated by the time and space complexity of the flow algorithm. Thus each flow computation either declares that the densest subgraph has density “at most $\lambda$”, or produces some $U\subseteq V$ with density strictly more than $\lambda$.
@@
Appendix E Public Bound on Arboricity
Our edge-DP matching algorithm (Theorem 4.1), node-DP matching algorithms (Theorem 4.13, Theorem 4.14), and node-DP implicit vertex cover algorithm (Theorem 5.6) all require a public data-oblivious estimate $\tilde{\alpha}$ of the maximum arboricity $\alpha$ over the stream. While our algorithms always guarantee privacy, their utility is only guaranteed when $\tilde{\alpha}\geq\alpha$. One way to remove this assumption is to first execute an insertion-only continual release algorithm for estimating the arboricity and then run the desired algorithm for a second pass given the estimated arboricity.
E.1 Notification of Failure
@@
Hence our $k$-core algorithm immediately yields a $(4+\eta,O(\varepsilon^{-1}\eta^{-2}\log^{3}(n)))$-approximation of the arboricity. When this estimated arboricity (approximately) exceeds the public bound, we can notify the observer that utility is no longer guaranteed by outputting “FAIL”. Note that algorithms with better approximation and space guarantees that directly estimate $\alpha$ most likely exist but we describe the approach here using Theorem 7.1 for the sake of simplicity.
E.2 Guessing the Arboricity
@@
In the case of our matching algorithms (Section 4), we briefly sketch how to completely remove the dependence on a prior public $\tilde{\alpha}$ bound at the cost of additional space. Suppose we are given an $\varepsilon$-edge DP $(\beta,\zeta)$-approximation continual release algorithm $\mathcal{A}_{e}$ for computing the arboricity that uses $O(S)$ space (e.g. see previous section) or similarly an $\varepsilon$-node DP approximation algorithm $\mathcal{A}_{n}$. We can guess $\tilde{\alpha}=2^{p}$ for each $p=1,2,\dots,O(\log n)$ and execute $O(\log n)$ instances of our edge- or node-DP continual release matching algorithm (Theorem 4.1, Theorem 4.13) alongside $\mathcal{A}$. Thus we can identify the “correct” output from the parallel instances corresponding to the current arboricity $\alpha_{t}$ up to a $O\left(2\beta,O(\zeta)\right)$-approximation.

### v3

@@ E.1 Inner Product Queries
Theorem E.1 (Theorem 4.5 in [jain2023counting]).
@@
E.2 Maximum Cardinality Matching
We now restate and prove the lower bound for computing the size of a maximum matching in the fully-dynamic continual release model. See 8.1
E.2.1 Description of Reduction
@@ 20       $r^{(0)},r^{(1)},\dots,r^{(k)}\leftarrow\mathcal{A}(S)$ answers to queries
Algorithm E.1 Reduction from Inner Product Queries to Maximum Matching
E.2.2 Proof of Lower Bound
Lemma E.2.
Given an $\varepsilon$-DP mechanism for maximum matching in the continual release setting that outputs estimates within additive error at most $\zeta$ with probability $0.99$ for a fully-dynamic stream of length $T\geq n+2nk$, Algorithm E.1 is an $\varepsilon$-DP mechanism for answering $k$ inner product queries within additive error $2\zeta$ with probability $0.99$.
Proof (Lemma E.2).
Let $\mathcal{M}$ be such a $\varepsilon$-DP mechanism and suppose we are given a private dataset $y\in\{0,1\}^{n}$ as well as public linear queries $q^{(1)},\dots,q^{(k)}$. Let $S^{(0)},S^{(1)},\dots,S^{(k)}$ be the partial update streams as in Algorithm E.1 and
@@ E.2.2 Proof of Lower Bound
Proof (Theorem E.3).
Let $c_{1},c_{2}$ be the constants from Theorem E.1. Fix $n\in\mathbb{N}$ and suppose $T\geq n+2c_{1}n^{2}$. By Lemma E.2, $\mathcal{A}$ implies a $1$-DP mechanism for answering $k=c_{1}n$ inner product queries within additive error $2\zeta$ with probability $0.99$. But then by Theorem E.1, we must have $\zeta\geq c_{2}\sqrt{n}/2$.
@@
In the regime $T\geq\nicefrac{{1}}{{\varepsilon}}$, combining Lemma E.4 and Theorem E.3 implies a lower bound of
@@
E.3 Connected Components
We now prove Theorem 8.2, whose statement we repeat below for convenience. See 8.2
E.3.1 Description of Reduction
@@ 24             Output $\lVert q^{(j)}\rVert_{0}+r^{{(j)}}_{n}-r^{(0)}_{3n}$
Algorithm E.2 Reduction from Inner Product Queries to Connected Components
E.3.2 Proof of Lower Bound
Lemma E.5.
Given an $\varepsilon$-DP mechanism for connected components in the continual release setting that outputs estimates within additive error at most $\zeta$ with probability $0.99$ for a fully-dynamic stream of length $T\geq 3n+2nk$, Algorithm E.2 is an $\varepsilon$-DP mechanism for answering $k$ inner product queries within additive error $2\zeta$ with probability $0.99$.
Proof (Lemma E.5).
Let $\mathcal{M}$ be such a $\varepsilon$-DP mechanism and suppose we are given a private dataset $y\in\{0,1\}^{n}$ as well as public linear queries $q^{(1)},\dots,q^{(k)}$. Let $S^{(0)},S^{(1)},\dots,S^{(k)}$ be the partial update streams as in Algorithm E.2 and
@@ E.3.2 Proof of Lower Bound
Proof (Theorem E.6).
Let $c_{1},c_{2}$ be the constants from Theorem E.1. Fix $n\in\mathbb{N}$ and suppose $T\geq 3n+2c_{1}n^{2}$. By Lemma E.5, $\mathcal{A}$ implies a $1$-DP mechanism for answering $k=c_{1}n$ inner product queries within additive error $2\zeta$ with probability $0.99$. But then by Theorem E.1, we must have $\zeta\geq c_{2}\sqrt{n}/2$.
@@
By following an identical argument to Lemma E.4, i.e. creating duplicate instances of the lower bound construction within a longer stream, we derive a similar lower bound to Theorem 8.1 that accounts for the privacy parameter $\varepsilon$.
Appendix F Private Static Densest Subgraph
Theorem F.1 (Theorem 6.1 in [DLL23]).
@@ Appendix F Private Static Densest Subgraph
Theorem F.2 (Theorem 5.1 in [DLRSSY22]).
Fix $\eta\in(0,1]$. There is an $\varepsilon$-edge DP densest subgraph algorithm that runs in $O((m+n)\log^{3}n)$ time and $O(m+n)$ space and returns a subset $V^{\prime}\subseteq V$ of vertices that induces a $O\left(1+\eta,O(\varepsilon^{-1}{\eta^{-3}}\log^{4}n)\right)$-approximate densest subgraph with probability at least $1-O(1/\operatorname*{poly}(n))$ for any constant $\eta>0$.
Appendix G The Density of the Densest Subgraph
@@
Proposition G.2 ([chekuri2022densest]).
@@
[…]
Theorem G.3 ([dinitz2006dinitz]).
@@
We are now ready to show Theorem G.1.
Proof (Theorem G.1).
@@
Construct the auxiliary flow network $D=(N,A)$ above for some guess $\lambda\geq\nicefrac{{\lvert E\rvert}}{{\lvert V\rvert}}$ of the value of the densest subgraph. There are $2+\lvert V\rvert+\lvert E\rvert=O(\lvert E\rvert)$ nodes and $\lvert E\rvert+2\lvert E\rvert+\lvert V\rvert=O(\lvert E\rvert)$ arcs. We can check if the densest subgraph has value at most $\lambda$ by computing the maximum flow in the $D$ using Dinic’s algorithm Theorem G.3. This is justified by Proposition G.2. Moreover, each flow computation also retrieves the vertices inducing a minimum $(s,t)$-cut by computing the vertices reachable from $s$ in the residual graph. The time and space complexity of computing the residual flow is dominated by the time and space complexity of the flow algorithm. Thus each flow computation either declares that the densest subgraph has density “at most $\lambda$”, or produces some $U\subseteq V$ with density strictly more than $\lambda$.
@@
Appendix H Public Bound on Arboricity
Our edge-DP matching algorithm (Theorem 7.1), requires a public data-oblivious estimate $\tilde{\alpha}$ of the maximum arboricity $\alpha$ over the stream. While our algorithms always guarantee privacy, their utility is only guaranteed when $\tilde{\alpha}\geq\alpha$. One way to remove this assumption is to first execute an insertion-only continual release algorithm for estimating the arboricity and then run the desired algorithm for a second pass given the estimated arboricity.
H.1 Notification of Failure
@@
Hence our $k$-core algorithm immediately yields a $(5,O(\varepsilon^{-1}\log^{3}(n)))$-approximation of the arboricity. When this estimated arboricity (approximately) exceeds the public bound, we can notify the observer that utility is no longer guaranteed by outputting “FAIL”. Note that algorithms with better approximation and space guarantees that directly estimate $\alpha$ most likely exist but we describe the approach here using Theorem 5.1 for the sake of simplicity.
H.2 Guessing the Arboricity
@@
In the case of our matching algorithms (Section 7), we briefly sketch how to completely remove the dependence on a prior public $\tilde{\alpha}$ bound at the cost of additional space. Suppose we are given an $\varepsilon$-edge DP $(\beta,\zeta)$-approximation continual release algorithm $\mathcal{A}_{e}$ for computing the arboricity that uses $O(S)$ space (e.g. see previous section). We can guess $\tilde{\alpha}=2^{p}$ for each $p=1,2,\dots,O(\log n)$ and execute $O(\log n)$ instances of our edge-DP continual release matching algorithm (Theorem 7.1) alongside $\mathcal{A}$. Thus we can identify the “correct” output from the parallel instances corresponding to the current arboricity $\alpha_{t}$ up to a $O\left(2\beta,O(\zeta)\right)$-approximation.